# Diederich–Fornæss index and global regularity of the complex Green operator: domains with comparable Levi eigenvalues

### v1

@@ 6. Appendix: Kohn–Morrey–Hörmander on the boundary
(6.1) $\|\bar{\partial}_{b}u\|_{\varphi}^{2}+\|\bar{\partial}_{b,\varphi}^{\star}u\|_{\varphi}^{2}\gtrsim\;\|\overline{L}u\|_{\varphi}^{2}+2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\left(T-\frac{1}{2}\frac{\partial\varphi}{\partial\nu}\right)u_{jK}\overline{u_{kK}}e^{-\varphi}\\ +2\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}\varphi}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}e^{-\varphi}+\mathcal{O}(\|u\|_{\varphi}^{2}),$
where $\|\cdot\|_{\varphi}$ is the weighted norm with respect to the weight function $e^{-\varphi}$, $\|\overline{L}u\|_{\varphi}^{2}$ is the sum $\sum_{j}\sum_{J}^{\prime}\|\overline{L}_{j}u_{J}\|_{\varphi}^{2}$, $K$ is a multi-index of length q1q-1, $T=L_{n}-\overline{L}_{n}$, $c_{jk}$ are the coefficients of the Levi form in the basis $\{L_{1},\dots,L_{n-1}\}$, $u=\sum_{J}^{\prime}\widetilde{u_{J}}d\bar{z}_{J}$ is the local expression of $u$ in the Euclidean coordinates, and the constant appearing in the $\mathcal{O}$ term is independent of $\varphi$.
Estimate (6.1) is a special case of an estimate in [19]. We thank Andy Raich for pointing us to the second formula from the bottom on page 150. Lets call it (HR). In (HR), take $\lambda^{+}=\varphi$, $m=0$ (since $\Omega$ is pseudoconvex), and $\phi=u$. Then the left hand side of (HR) gives the left hand side of (6.1), the first sum in the fist line of (HR) is absent, and the second sum equals the first term on the right hand side of (6.1). The coefficients $h^{+}_{jk}$ are defined in Lemma 4 in [19] as $h^{+}_{jk}=c_{jk}-\delta_{jk}\frac{1}{q}\sum_{l=1}^{m}c_{ll}=c_{jk}$ (because $m=0$), where $c_{jk}$ are the coefficients of the Levi form; i.e. they are the same as in (6.1). Next, $\Gamma^{\lambda^{+}}_{jk}=\Theta^{\lambda^{+}}_{jk}$ by the formula immediately preceding (HR) (since $m=0$). Definition (1) of $\Theta^{\lambda^{+}}$ and Proposition 3.1 now show that $\sum_{I\in\mathcal{J}_{q-1}}\sum_{j,k=1}^{n-1}\text{Re}(\Gamma^{\lambda^{+}}_{jk}\phi_{jI},\phi_{kI})=\sum^{{}^{\prime}}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}\varphi}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}e^{-\varphi}$. Therefore, (HR) translates to
(6.2) $\|\bar{\partial}_{b}u\|_{\varphi}^{2}+\|\bar{\partial}_{b,\varphi}^{\star}u\|_{\varphi}^{2}=\|\overline{L}u\|_{\varphi}^{2}+\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\left(T-\frac{1}{2}\frac{\partial\varphi}{\partial\nu}\right)u_{jK}\overline{u_{kK}}e^{-\varphi}\\ +\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}\varphi}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}e^{-\varphi}+E(u)\;,$
@@
(6.3) $\|\bar{\partial}_{b}u\|_{\varphi}^{2}+\|\bar{\partial}_{b,\varphi}^{\star}u\|_{\varphi}^{2}\gtrsim\frac{1}{2}\|\overline{L}u\|_{\varphi}^{2}+\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\left(T-\frac{1}{2}\frac{\partial\varphi}{\partial\nu}\right)u_{jK}\overline{u_{kK}}e^{-\varphi}\\ +\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}\varphi}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}e^{-\varphi}+\mathcal{O}(\|u\|_{\varphi}^{2})\;$;

### v2

@@ 6. Appendix: Kohn–Morrey–Hörmander on the boundary
(6.1) $\|\bar{\partial}_{b}u\|_{\varphi}^{2}+\|\bar{\partial}_{b,\varphi}^{\star}u\|_{\varphi}^{2}\gtrsim\;\|\overline{L}u\|_{\varphi}^{2}+2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\left(T-\frac{1}{2}\frac{\partial\varphi}{\partial\nu}\right)u_{jK}\overline{u_{kK}}e^{-\varphi}\\ +2\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}(\partial\overline{\partial}\varphi)\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)e^{-\varphi}+\mathcal{O}(\|u\|_{\varphi}^{2}),$
where $\|\cdot\|_{\varphi}$ is the weighted norm with respect to the weight function $e^{-\varphi}$, $\|\overline{L}u\|_{\varphi}^{2}$ is the sum $\sum_{j}\sum_{J}^{\prime}\|\overline{L}_{j}u_{J}\|_{\varphi}^{2}$, $K$ is a multi-index of length q1q-1, $T=L_{n}-\overline{L}_{n}$, $c_{jk}$ are the coefficients of the Levi form in the basis $\{L_{1},\dots,L_{n-1}\}$, $\tilde{L}_{u}^{K}$ are the vector fields from section 3, and the constant appearing in the $\mathcal{O}$ term is independent of $\varphi$.
Estimate (6.1) is a special case of an estimate in [19]. We thank Andy Raich for pointing us to the second formula from the bottom on page 150. Lets call it (HR). In (HR), take $\lambda^{+}=\varphi$, $m=0$ (since $\Omega$ is pseudoconvex), and $\phi=u$. Then the left hand side of (HR) gives the left hand side of (6.1), the first sum in the fist line of (HR) is absent, and the second sum equals the first term on the right hand side of (6.1). The coefficients $h^{+}_{jk}$ are defined in Lemma 4 in [19] as $h^{+}_{jk}=c_{jk}-\delta_{jk}\frac{1}{q}\sum_{l=1}^{m}c_{ll}=c_{jk}$ (because $m=0$), where $c_{jk}$ are the coefficients of the Levi form; i.e. they are the same as in (6.1). Next, $\Gamma^{\lambda^{+}}_{jk}=\Theta^{\lambda^{+}}_{jk}$ by the formula immediately preceding (HR) (since $m=0$). Definition (1) of $\Theta^{\lambda^{+}}$ and Proposition 3.1 now show that $\sum_{I\in\mathcal{J}_{q-1}}\sum_{j,k=1}^{n-1}\text{Re}(\Gamma^{\lambda^{+}}_{jk}\phi_{jI},\phi_{kI})=\sum^{{}^{\prime}}_{K}(\partial\overline{\partial}\varphi)\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)$. Therefore, (HR) translates to
(6.2) $\|\bar{\partial}_{b}u\|_{\varphi}^{2}+\|\bar{\partial}_{b,\varphi}^{\star}u\|_{\varphi}^{2}=\|\overline{L}u\|_{\varphi}^{2}+\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\left(T-\frac{1}{2}\frac{\partial\varphi}{\partial\nu}\right)u_{jK}\overline{u_{kK}}e^{-\varphi}\\ +\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}(\partial\overline{\partial}\varphi)\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)e^{-\varphi}+E(u)\;,$
@@
(6.3) $\|\bar{\partial}_{b}u\|_{\varphi}^{2}+\|\bar{\partial}_{b,\varphi}^{\star}u\|_{\varphi}^{2}\gtrsim\frac{1}{2}\|\overline{L}u\|_{\varphi}^{2}+\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\left(T-\frac{1}{2}\frac{\partial\varphi}{\partial\nu}\right)u_{jK}\overline{u_{kK}}e^{-\varphi}\\ +\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}(\partial\overline{\partial}\varphi)\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)e^{-\varphi}+\mathcal{O}(\|u\|_{\varphi}^{2})\;$;